Suppose a recurrence relation




an=2an−1−an−2


where a1=7 and a2=10




can be represented in explicit formula, either as:


Formula 1:


an=pxn+qnxn


       or  


Formula 2:


an=pxn+qyn


 


where 


x


and


y


are roots of the characteristic equation.




  Determine p and q




1
Expert's answer
2021-11-30T17:43:45-0500

Let us solve the characteristic equation of the recurrence relation "a_n=2a_{n\u22121}\u2212a_{n\u22122}," which is equivalent to "a_n-2a_{n\u22121}+a_{n\u22122}=0." It follows that the characteristic equation "x^2-2x+1=0" is equivalent to "(x-1)^2=0," and hence has the roots "x_1=x_2=1." It follows that the solution of the recurrence equation is "a_n=p\cdot 1^n+q\cdot n1^n=p+q\cdot n." Since "a_1=7" and "a_2=10," we conclude that "7=a_1=p+q" and "10=a_2=p+2q." Therefore, "q=3" and "p=4."

Answer: "p=4,\ q=3."